Find the length of the leg of this right triangle. Give an approximation to 3 decimal places.
the triangles sides are 35, 30, and a


a. 8.062 c. 46.098
b. 17.748 d. 18.028

Answer by rfer(16322)

35^2+30^2=c^2
1225+900=2125
c=46.098
C)
It could be answer D)18.028
It says leg, so the way it is writen it is hard to tell
if they are looking for the diagonal or one of the legs.